OpenStudy (anonymous):

Admission to the zoo is $6 for children and $9 for adults. On a certain day, 1480 people go to the zoo and $10,500 is collected. How many children and how many adults went to the zoo?

OpenStudy (dan815):

X+Y=1480 6X+9Y=10500

OpenStudy (anonymous):

i need how many were chidrin and how many were adults please

OpenStudy (dan815):

SOLVE THE EQN

OpenStudy (anonymous):

do you know what the X and Y are @dan815 i need them i not know how to get them

OpenStudy (texaschic101):

a + c = 1480 ----- adults + children = 1480 9a + 6c = 10,500 --- $9 per adult ticket + $6 per child ticket = 10,500 These are the equations you will need to use. We can solve this by substitution. a + c = 1480 -- rearrange a = 1480 - c now sub 1480 - c in for a in the 2nd equation 9a + 6c = 10500 9(1480 - c) + 6c = 10500 --- distribute through the parenthesis 13320 - 9c + 6c = 10500 --- subtract 13320 from both sides -9c + 6c = 10500 - 13320 -- combine like terms -3c = - 2820 --- divide both sides by -3 c = -2820/-3 c = 940 now sub 940 in for c in either equation to find a a + c = 1480 a + 940 = 1480 a = 1480 - 940 a = 540 check.. 9a + 6c = 10500 9(540) + 6(940) = 10500 4860 + 5640 = 10500 10500 = 10500 (correct) so adults present (a) = 540, and children present (c) = 940 any questions at all ?